Figure 1A. U.S. Output and the Measured Labor Wedge

Annual 1929–39, Series Normalized to Equal 100 in 1929

Source: Chari, Kehoe, and McGrattan (2007)

Figure 1B. U.S. Labor and Prediction of Model

with Just the Labor Wedge

Annual 1929–39, Series Normalized to Equal 100 in 1929

Source: Chari, Kehoe, and McGrattan (2007)

Figure 2A. U.S. Output and Prediction of Smets-Wouters (2007) Model with the Dubiously Structural Shocks^{∗}

Quarterly Percentage Changes, 1965–2005, Series Logged and Detrended

∗The dubiously structural shocks include the wage-markup shock, the price-markup shock, the exogenous spending shock, and the risk premium shock.

Figure 2B. U.S. Hours and Prediction of Smets-Wouters (2007) Model with the Dubiously Structural Shocks^{∗}

Quarterly Percentage Changes, 1965–2005, Series Logged and Demeaned

∗The dubiously structural shocks include the wage-markup shock, the price-markup shock, the exogenous spending shock, and the risk premium shock.

Figure 2C. U.S. Inflation and Prediction of Smets-Wouters

(2007) Model with the Dubiously Structural Shocks^{∗}

Quarterly, 1965–2005, Series Demeaned

∗The dubiously structural shocks include the wage-markup shock, the price-markup shock, the exogenous spending shock, and the risk premium shock.

Figure 3. U.S. Output and Potential Output in Version

of Smets-Wouters (2007) Model with AR(1) Taste

Shocks and i.i.d. Wage-Markup Shocks

Quarterly Percentage Changes, 1965–2005, Series Logged and Detrended

Figure 4. Annualized Interest Rate and Risk Premium Shock

of the Smets-Wouters (2007) Model

Quarterly 1965–2005, Series Demeaned

Figure 5. Price of Angel Soft Bathroom Tissue at Chicago’s Dominick’s

Finer Foods and Price Implied by Backward Indexation

Weekly, from Week 11 of 1991 to Week 5 of 1993

Source of Actual Data: University of Chicago, Kilts Center for Marketing